judy frog hops up a 2/3rd of the ladder, slides down a 1/7th of the ladder from where she is at to the ground and then climbs up 2/5th of the remaining ladder distance to the top. How far is judy off the ground?
Answer by greenestamps(13203)



Let the height of the ladder be x.

judy frog hops up a 2/3rd of the ladder...

She is now at a height equal to 2/3 of the ladder: (2/3)x

... slides down a 1/7th of the ladder from where she is at to the ground...

The distance she slides down is 1/7 of (2/3)x; the distance sheis now above the ground is 6/7 of (2/3)x, which is (12/21)x = (4/7)x

... and then climbs up 2/5th of the remaining ladder distance to the top.

She is at 4/7 the height of the ladder; the distance remaining to the top is 3/7 of the height of the ladder. She climbs 2/5 of that distance:

(2/5)*(3/7)x = (6/35)x

The distance she is now off the ground is (4/7)x + (6/35)x = (20/25)x + (6/35)x = (26/35)x.

ANSWER: She finishes 26/35 of the height of the ladder off the ground.